# How Quantum Complexity Helps Classical Complexity: A Survey

### 16 Dec 2023 v1

@@ 3.3. Approximation for Quantum Problems
The Tail Assignment problem is the task of assigning individual aircraft to a given set of flights, minimizing the overall cost. In this article, Vikstål et al. (vikstaal2020applying, ) simulated the QAOA applied to instances of this problem derived from real-world data of the Tail Assignment problem. These instances were reduced to fit on quantum devices with 8, 15, and 25 qubits. For instance, the reduction process leaves only one viable solution that allows us to map out an Exact Cover problem on Tail Assignment (vikstaal2020applying, ).
@@ 4.2. Maximum Independent Set
In the above definitions, $V$ is the set of vertices of which a component $(j,k)$ corresponds to the literal $l_{j,k}$ in the $j$-th clause $C_{j}$. Also, $E$ is the set of all edges, namely the union of two edge sets $E_{1}$ and $E_{2}$. In particular, $E_{1}$ connects vertices in the same clause and $E_{2}$ connects vertices in different clauses negating each other.
@@
There are two limitations to the Rydberg atoms system as a quantum simulator. First, one may not simulate non-planar graphs with 2D Rydberg atoms. Second, graphs with high-degree vertices cannot be encoded. This is because the blockade radius sets the size of Rydberg atom interactions. Consider the quantum wires as chains of auxiliary wire atoms that mediate strong interactions between distant atoms. Using quantum wires, one can tackle MIS (kim2022rydberg, ).
Finding an independent set with the largest possible size for a given graph is required to solve the maximum independent set problem. It is known that this problem is NP-hard, which means that there is unlikely to be an efficient algorithm to find the maximum independent set of graphs. This article gives us the first results of experiments involving arbitrarily structured MIS inputs using a D-Wave QPU (Commercial quantum processing unit), which is increasingly used to solve complex combinatorial optimization problems. The approach of Yarkoni et al. (8477865, ) shows how one may map MIS to QUBO problems. In their work, the authors had a limiting factor, the number of QPU (Quantum Processing Unit) parameters to be tuned to optimize performance (8477865, ).
@@ 4.3. Graph Traversal Problems
Warren (warren2013adapting, ) shows how to guide a quantum computer in selecting an optimal tour for the traveling salesman. Adiabatic quantum computers are excellent machines for solving TSPs. They solve discrete optimization problems and have computational capacity and speed attributed to their quick selection of a near-optimal outcome from an exponential number of possibilities. There are currently limited numbers of binary variables that can be handled by adiabatic quantum computers. This is because, in the Ising model, binary variables are mixed with physical quantum bits. According to this article, as the number of quantum bits grows, it is no longer possible to build a connection between them through existing practical development methods. This means that the size of TSPs that can be implemented is limited (warren2013adapting, ).
@@ 5.1. QAOA
It is known that the gap between the QAOA and the abQAOA (adaptive-bias QAOA) is greatest when the plateau barrens  [^3] The barren plateau phenomenon explains that the gradient of parameters of the parameterized quantum circuits will vanish exponentially in terms of the system size (Zhao_2021, ) are most visible. Equally, numerical evidence suggests that abQAOA has better-scaling behavior than QAOA in terms of system size accuracy (See (yu2023solution, ) for more information). Is it possible to study a barren plateau phenomenon with abQAOA?
@@ 5.4. Quantum Walk
An algorithm is suggested in the work in (laarhoven2016search, ) for quantum random walks, in which there are no better generic algorithms for finding $k$ different marks than to run the whole random walk (including the setup) $\mathcal{O}(k)$ times. Is it possible to do this more effectively and improve the overall complexity of the algorithm?
@@
Can an enhancement to the Shortest Vector Problem (SVP), as outlined in the work by Chailloux et al. (chailloux2021lattice, ) in their paper Lattice-Based Cryptography, involve incorporating a layer of ”local sensitive filtering” through a quantum random walk, potentially by introducing a local sensitivity value within the random walk graph, as an alternative to using the Johnson graph?
@@ 5.5. Mean Estimator
. Nonetheless, the quantum settings in (cornelissen2022near, ) yield
@@
. As a future direction, one can investigate whether some combination of the above approaches may be proved optimal under all settings.
@@ 5.8. Quantum String Matching Algorithm
It is known that String Synchronizing Sets has applications in classical string algorithms. Are there more applications in quantum string algorithms? For E.g. given a maximum of $k$, what is the quantum query complexity to compute the edit distance of two strings s.t. their distance is no larger than $k$?
@@ 5.9. Quantum Merging Trees
[…]
@@
Is it possible to extend the merging trees (naya2020optimal, ), inspired by the classical literature on k-list problems? Some of such techniques have been explored that did not result in a quantum advantage.
@@ 5.11. Verification and Certification
The work in (zhang2022classical, ) pointed out that the number of required rounds for its remote state preparation protocol runs is constant. Nonetheless, the number of rounds for its gadget-assisted verification protocol is proportional to linear. So, the overall number of rounds for its classical verification of quantum computation protocol is linear. Is it possible to overcome this limitation?
@@
It seems challenging to design a sampling-based quantum supremacy experiment that runs on a NISQ device and admits classical verification. Is it possible to repurpose the results discussed in (aaronson2023certified, ) and obtain a secure and verifiable certified randomness scheme that runs on existing devices?
@@ 5.13. Complexity
Consider bounding the time complexity of the guessing tree method for solving the maximum matching problem. Such a result allows comparing the maximum matching algorithm described in (kimmel2021query, ) to existing quantum maximum matching algorithms that aim to minimize time complexity rather than query complexity.
@@
Consider repeated applications of Grover’s search (dorn2009quantum, ). Is it possible to achieve a time complexity within a logarithmic factor of the query complexity that appeared in the work discussed in (witter2022query, )?
@@
For maximum matching in general graphs, there is a query complexity of $\mathcal{O}(\sqrt{m}n^{\frac{3}{4}})$ (witter2022query, ). Is it possible to achieve a better time complexity through repeated applications of Grover’s search, within a logarithmic factor of the aforementioned query complexity?
@@
Using a problem inspired by Brakerski et al.’s proof of quantumness method allows for proving more fine-grained separations between hybrid classes. For instance
@@
. One might then ask whether these separations yield finer-grained proofs of quantum depth as well. This cannot be immediately concluded from the existing work (for instance, see (arora2023quantum, )).
@@
It is already known that there exist families of Hamiltonians with all low energy states of non-trivial complexity (with complexity measured by the quantum circuit depth preparing the state). Is it possible to generalize the proof techniques of the latest results (See (anshu2023nlts, )) to prove non-trivial lower bounds for non-commuting Hamiltonians?
@@ 5.14. Classification
It is known that non-directed graphs can be encoded with $n$ nodes using quantum gates on $n$ qubits. The relation between the adjacency characteristics of the graph nodes may be set up to the angles of each amplitude of the system’s quantum state. An existing algorithm deals with an exponential number of graphs within polynomial time. It also provides insight into invariant(s) on graph isomorphism (de2022quantum, ). But under what circumstances non-isomorphic-graphs can share the same classification table of subgraphs by the number of edges? Also, how may the aforementioned algorithm classify the graphs considered hard to classify by classical algorithms e.g. strong-regular-graphs?
@@
Mycielski transformation is a method used in graph theory to construct a new graph from an existing one. The Mycielski transformation is also defined for quantum graphs and affects (quantum) clique numbers and (quantum) chromatic numbers. The Motzkin-Straus clique number is a way of finding the largest possible clique in a given graph. It is based on a formula that involves the eigenvalues of the adjacency matrix of the graph. Characterize cones that correspond to the clique numbers for quantum graphs. Which Motzkin-Straus clique numbers does the Mycielski transformation preserve?
@@ 5.16. Others
[…]
Consider the Quantum-Accessible Quantum Memory (QAQM) model in quantum computation, where the quantum computation can use as many qubits as needed. Are there better algorithms for the above problem without QAQM or even in the low-qubits model?
@@
The optimization problem of maximum cardinality matching has already been applied to the latest QA hardware from D-Wave, the 2000Q. This involved an improved version of the QUBO formulation of this problem. Fabricated graphs to the size of $22$ variables (edges) have been used in their energy gap scale investigations (mcleod2022benchmarking, ). Is it possible to extend this analysis to a $G_{2}$ graph? One will then gain direct insight into the original Graph family and be able to verify its scaling behavior with the maximum match problem for the $G_{n}$ graph family.
@@
Consider the Measurement-Based Quantum Computation (MBQC) model. Transforming a circuit in the usual circuit model suggests yet another blowup for the circuit width. One may apply the usual circuit model gates onto arbitrary wires. Nonetheless, only the interactions between neighbors are supported in the MBQC model. In addition, using long-range gates is costly. Is there a way to suggest a protocol for both models? Also, one may focus on a protocol for the usual circuit model.
@@
The construction of the work in (anshu2023nlts, ) requires no quantum local testability. The property proved in the above work (Property 1) suffices to cluster the classical distributions of low-energy states. But note that it is weaker than local testability. It is already discussed in the literature that local testability implies NLTS. What codes with Property 1 may imply for the quantum PCP conjecture (aharonov2013guest, )?
@@
Consider the property proved in the work discussed in (anshu2023nlts, ), referred to as Property 1. From the chain complexes perspective, Property 1 appears to be related to the small-set boundary and co-boundary expansion in Definition 1.2 of (hopkins2022explicit, ). Does this introduce a classical analog of the NLTS property? There is a construction in (hopkins2022explicit, ) of classical Hamiltonians, hard to approximate, made by using the sum-of-squares hierarchy.
@@
An existing method solves the MIS problem using quantum wires for the case of non-planar graphs and graphs with high dimensions. Quantum wires are a set of auxiliary wire atoms and, to synthesize an entire target graph, $G_{T}$, out of the simple first graph, $G_{0}$, they have been proposed to mediate strong interactions between distant atoms (kim2022rydberg, ). Can the quantum wire be used for other optimization problems than the maximum independent set?
@@
It is known that both the minimum weight and the maximum weight matching problems can be solved in time $\mathcal{O}(n(m+n\log n))$. Also, an algorithm already exists in which a maximum weight augmenting path in graph $G$ can be found in time $\mathcal{O}(m+n\log n)$. The maximum weight matching will be found with $n$ steps after enlarging in $G$ the matching with the path. Is there a quantum algorithm for the minimum weight and the maximum weight matching for general graphs? This algorithm may improve the running time of the classical algorithm (dorn2009quantum, ).
@@
A method already exists in the literature for generating arbitrarily structured maximum independent set problems and preparing them for submission to D-Wave quantum processing units (QPU) (8477865, ). Can iterative parameter tuning and finding additional problem classes (like the MIS) be incorporated, which are both classically difficult and low precision in the Quadratic unconstrained binary optimization (QUBO) form? This makes it easier to compare QPU performance with classical algorithms and gives an overview of use cases for quantum annealing in real applications.

### 10 Jan 2024 v2

# Where Quantum Complexity Helps Classical Complexity: A Survey

@@ 2. Summary
2.1. Statistical Analysis
Table 1 presents a comparison of time complexities for various problems in the context of Quantum versus Classical computing. The table consists of two columns: ”Classic C.” and ”Quantum C.”, each providing information based on the latest level of our knowledge in the field.
There is an additional column labeled ”Improvement.” This column aims to clarify whether a Quantum solution can be considered an improvement for a given problem. If a quantum computer solves a problem in time $T$ and a classical computer requires a polynomial function of $T$ (such as $T^{2}$), it is termed a polynomial speedup. When the time complexity of the best classical algorithm, denoted as $C(n)$, and the time complexity of the quantum algorithm, denoted as $Q(n)$, satisfies $C=2^{\Omega(Q^{\lambda})}$ for some positive constant $\lambda$, the speedup is referred to as superpolynomial. Exponential speedup occurs when a quantum computer requires time $T$, while a classical computer needs an exponentially growing function of $T$ (for example, $2^{T}$). These categories represent the promising and desired speedups in quantum computers.
By Problem Number(s) in the last column, we refer to the corresponding open problem(s) mentioned in the paper. In this table, $|V|=n$ is the number of vertices and $|E|$ is the number of edges in a graph.
Problem Classic C. Quantum C. Improvement Problem Number(s) k-distinctness $\Theta(n\log n)$ $\tilde{\mathcal{O}}(n^{\frac{3}{4}-\frac{1}{4}\frac{1}{2^{k}-1}})$ Polynomial - (Element Distinctness) (In decision tree model (10.1145/800061.808735, ))  (jeffery2022multidimensional, ) k-SAT ($k\geq 3$) $\mathcal{O}(2^{1-\frac{\Theta(1)}{k}})$ (10.1145/3313276.3316359, ) $\mathcal{O}((2(1-1/k))^{n})$ (schoning1999probabilistic, ) - 1, 2, 15, 16, 17, 66 32, 33, 34, 35, 36 Maximum Independent Set $\mathcal{O}(1.1996^{n}\cdot n^{\mathcal{O}(1)})$ (XIAO2017126, ) $\mathcal{O}(1.1488^{n})$ (etde_21241168, ) Polynomial 67, 98, 104, 98 Hamiltonian Cycle $\mathcal{O}^{*}(1.657^{n})$ (bjorklund2014determinant, ) $\mathcal{O}(n^{2n/(n+1)})$ (dorn2007quantum, ) Exponential 18, 10, 11, 14 Eulerian Tour $\mathcal{O}(|E|)$ (fleischner1990eulerian, ) $\mathcal{O}(\sqrt{n})$ (dorn2007quantum, ) Polynomial - Travelling Salesman Problem $\mathcal{O}(n^{2}\cdot 2^{n})$ (doi:10.1137/0110015, ) $\mathcal{O}({1.728}^{n}\cdot\text{poly}(n))$ (doi:10.1137/1.9781611975482.107, ) - 14 Maximum Matching $\mathcal{O}(|E|\sqrt{|V|})$ (micali1980v, ) $n^{3/2}\log^{2}n$ (beigi2022time, ) Polynomial 68, 86, 69, 71, 103 Maximum Cut $\text{exp}(n)$ (garey1997computers, ) $\sqrt{2^{n}/r}$ (chang2023quantum, ) - 8, 6 s.t. $r$ is num of max-cuts Graph Isomorphism $2^{\mathcal{O}(\log n)^{c}}$ (grohe2020graph, ) NA (alagic2007quantum, ) NA 96, 79, 31, 58 Maximum Clique $2^{n/4}$ (robson2001finding, ) $\mathcal{O}(\log_{2}n\cdot\sqrt{2^{n}})$ (haverly2021comparison, ) - 88, 89 String Edit Distance $\mathcal{O}(s^{2}+n)$ (landau1998incremental, ) $\mathcal{O}(n^{2}/\log n)$ (equi2021quantum, ) Poly-logarithmic 74, 91, 92 s.t. $s$ is a max. edit distance Longest Common Subsequence $\mathcal{O}(n\log^{2}n)$ (bhowmick2019approach, ) $\tilde{\mathcal{O}}(\sqrt{n})$ (le2023quantum, ) Polynomial 42 Protein Folding $\omega(n^{c}),\ \forall c>0$ $\mathcal{O}(2^{n/2})$ (wong2022fast, ) - 27, 101, 102 Mixed-Integer Programming $(\log n)^{{}^{\mathcal{O}(n)}}\cdot\text{poly}(n)$ (reis2023subspace, ) $\exp(n)$ 13 Knapsack Problem $\mathcal{O}(2^{n/2})$ (doi:10.1137/0210033, ) $\mathcal{O}(2^{n/2})$ (fujimura2010quantum, ) - 39, 38, 5.12, 41 Subset-Sum $\tilde{\mathcal{O}}(2^{0.291n})$ (becker2011improved, ) $\tilde{\mathcal{O}}(2^{0.236n})$ (bonnetain2020improved, ) - 37, 44, 45, 46 Discrete Logarithm $\omega(n^{c}),\ \forall c>0$ $\mathcal{O}(\log|\mathcal{G}|)$ (hhan2023quantum, ) Superpolynomial - for a cyclic group $\mathcal{G}$ of prime order Integer Factorization $\exp(1.9(\log n)^{\frac{1}{3}}(\log\log n)^{\frac{2}{3}})$ (buhler1993factoring, ) ${\mathcal{O}}(b^{3})$ (shor1999polynomial, ) Superpolynomial - $n$ is the number to factor s.t. $b$ is num of bits of $n$ Crypto. Certified Random Bits - $\exp(n)$ (aaronson2023certified, ) - 50, 55
Table 1. Quantum vs. Classical complexity comparison for some known problems in computer science.
[…]
Table 2. Applicability of Quantum $*{}^{*}~{}$p to the best level of our knowledge, there was not much work in the literature that used the above techniques to improve the complexity of a classical problem. Still, some works suggested that the techniques might bring enhancements in the future.
@@ 3.3. Approximation for Quantum Problems
The Tail Assignment problem is the task of assigning individual aircraft to a given set of flights, minimizing the overall cost. Vikstål et al. (vikstaal2020applying, ) simulated the QAOA applied to instances of this problem derived from real-world data of the Tail Assignment problem. These instances were reduced to fit on quantum devices with 8, 15, and 25 qubits. For instance, the reduction process leaves only one viable solution that allows us to map out an Exact Cover problem on Tail Assignment (vikstaal2020applying, ).
@@ 4.2. Maximum Independent Set
In the above definitions, $V$ is the set of vertices of which a component $(j,k)$ corresponds to the literal $l_{j,k}$ in the $j$-th clause $C_{j}$. Also, $E$ is the set of all edges, namely the union of two edge sets $E_{1}$ and $E_{2}$. In particular, $E_{1}$ connects vertices in the same clause and $E_{2}$ connects vertices in different clauses negating each other. Figure 1 is an instance of the MIS graph.
Figure 1. The figure is taken from (jeong2023quantum, ). The MIS graphs reduced from the 3-SAT instance, where vertices represent literals $(x_{1},\cdots,x_{6}\text{ and negations})$, solid edges intra-clause logics, and dashed edges the inter-clause logics (between literals and their negations).
@@
There are two limitations to the Rydberg atoms system as a quantum simulator. First, one may not simulate non-planar graphs with 2D Rydberg atoms. Second, graphs with high-degree vertices cannot be encoded. This is because the blockade radius sets the size of Rydberg atom interactions. Consider the quantum wires as chains of auxiliary wire atoms that mediate strong interactions between distant atoms. Using quantum wires, one can tackle MIS (kim2022rydberg, ) (Figure 2).
Figure 2. Taken from  (kim2022rydberg, )
a) $K_{3,3}$ Graph, b) experimental $K_{3,3}$ in 3D (blue circles) making by quantum wires. c) The experimental probability distribution of $K_{3,3}$ be the solution of MIS
Finding an independent set with the largest possible size for a given graph is required to solve the maximum independent set problem. It is known that this problem is NP-hard, which means that there is unlikely to be an efficient algorithm to find the maximum independent set of graphs. The authors provided the first results of experiments involving arbitrarily structured MIS inputs using a D-Wave QPU (Commercial quantum processing unit), which is increasingly used to solve complex combinatorial optimization problems. The approach of Yarkoni et al. (8477865, ) shows how one may map MIS to QUBO problems. In their work, the authors had a limiting factor, the number of QPU (Quantum Processing Unit) parameters to be tuned to optimize performance (8477865, ).
@@ 4.3. Graph Traversal Problems
Warren (warren2013adapting, ) shows how to guide a quantum computer in selecting an optimal tour for the traveling salesman. Adiabatic quantum computers are excellent machines for solving TSPs. They solve discrete optimization problems and have computational capacity and speed attributed to their quick selection of a near-optimal outcome from an exponential number of possibilities. There are currently limited numbers of binary variables that can be handled by adiabatic quantum computers. This is because, in the Ising model, binary variables are mixed with physical quantum bits. As the number of quantum bits grows, it is no longer possible to build a connection between them through existing practical development methods. This means that the size of TSPs that can be implemented is limited (warren2013adapting, ).
@@ 5.1. QAOA
It is known that the gap between the QAOA and the abQAOA (adaptive-bias QAOA) is greatest when the plateau barrens  [^3] The barren plateau phenomenon explains that the gradient of parameters of the parameterized quantum circuits will vanish exponentially in terms of the system size (Zhao_2021, ) are most visible. Equally, numerical evidence suggests that abQAOA has better scaling behavior than QAOA in terms of system size accuracy (See (yu2023solution, ) for more information). Is it possible to study a barren plateau phenomenon with abQAOA?
@@ 5.4. Quantum Walk
Laarhoven (laarhoven2016search, ) proposed an algorithm for quantum random walks to locate $k$ different marks. The algorithm suggests running the random walk, including the setup, $\mathcal{O}(k)$ times, to find the marks. However, it remains unclear whether this approach can be optimized to improve the overall complexity of the algorithm.
@@
Can an enhancement to the Shortest Vector Problem (SVP), as outlined in the work by Chailloux et al. (chailloux2021lattice, ) in their paper Lattice-Based Cryptography, involve incorporating a layer of local sensitive filtering through a quantum random walk, potentially by introducing a local sensitivity value within the random walk graph, as an alternative to using the Johnson graph?
@@ 5.5. Mean Estimator
Nonetheless, the quantum settings in (cornelissen2022near, ) yield
@@
As a future direction, one can investigate whether some combination of the above approaches may be proved optimal under all settings.
@@ 5.8. Quantum String Matching Algorithm
It is known that String Synchronizing Sets has applications in classical string algorithms. Are there more applications in quantum string algorithms? E.g. given a maximum of $k$, what is the quantum query complexity to compute the edit distance of two strings s.t. their distance is no larger than $k$?
@@ 5.9. Quantum Merging Trees
The existing results achieved optimal results for all merging trees by implementing a set of optimization strategies that have been meticulously developed and refined (naya2020optimal, ). However, it is worth exploring the possibility of extending this framework to improve quantum algorithms further. Notably, the complexity of the $r$-th encryption merge tree algorithms exceeds the time limit of $2^{0.3n}$, raising the question of whether this bound can be surpassed. As the team continues to investigate this topic, it is crucial to consider how this framework can be optimized to meet the evolving demands of the field.
@@
Is it feasible to expand upon the merging trees (naya2020optimal, ), which draw inspiration from the classical literature on k-list problems? Although various techniques have been explored, none have yielded quantum advantages.
@@ 5.11. Verification and Certification
The work in (zhang2022classical, ) highlights the number of rounds required for two distinct protocols. The remote state preparation protocol demands a constant number of rounds, whereas the gadget-assisted verification protocol sees a linear increase in the number of rounds. As a result, the overall number of rounds for the classical verification of quantum computation protocol is linear. It is imperative to explore whether there exists a viable solution to overcome this limitation.
@@
Designing a sampling-based quantum supremacy experiment that runs on a NISQ device and admits classical verification seems challenging. Is it possible to repurpose the results discussed in (aaronson2023certified, ) and obtain a secure and verifiable certified randomness scheme that runs on existing devices?
@@ 5.13. Complexity
It is suggested to limit the time complexity of the guessing tree method used for solving the maximum matching problem. This restriction would enable a comparison between the maximum matching algorithm explained in (kimmel2021query, ) and the current quantum maximum matching algorithms that prioritize minimizing time complexity over query complexity.
@@
In consideration of the repeated application of Grover’s search, as demonstrated in the study in (dorn2009quantum, ), the question arises as to whether it is possible to achieve a time complexity that is within a logarithmic factor of the query complexity observed in the investigation discussed in (witter2022query, ).
@@
The problem of maximum matching in general graphs has a query complexity of $\mathcal{O}(\sqrt{m}n^{\frac{3}{4}})$ (witter2022query, ). However, is it possible to improve the time complexity of this problem by employing repeated applications of Grover’s search, while ensuring a logarithmic factor reduction from the aforementioned query complexity?
@@
Using a problem inspired by Brakerski et al.’s proof of quantumness method (brakerski_et_al:LIPIcs:2020:12067, ) allows for proving more fine-grained separations between hybrid classes. For instance
@@
One might then ask whether these separations yield finer-grained proofs of quantum depth as well. This cannot be immediately concluded from the existing work (for instance, see (arora2023quantum, )).
@@
The existence of Hamiltonians with low energy states that exhibit non-trivial complexity, as determined by the quantum circuit depth required to prepare them, is well-established. However, the question of whether the proof techniques employed in the latest research (as outlined in (anshu2023nlts, )) can be extended to yield non-trivial lower bounds for non-commuting Hamiltonians remains open.
@@ 5.14. Classification
Quantum gates on $n$ qubits can encode non-directed graphs with $n$ nodes. An algorithm can handle an exponential number of graphs within polynomial time and provides insight into invariants on graph isomorphism (de2022quantum, ). However, it is unclear under what circumstances non-isomorphic graphs can share the same classification table of subgraphs by the number of edges. Moreover, the algorithm’s efficacy in classifying graphs that are considered hard to classify by classical algorithms, such as strong-regular graphs, remains unknown.
@@
Mycielski transformation is a graph theory method that also applies to quantum graphs. It affects clique and chromatic numbers. Motzkin-Straus clique number finds the largest possible clique using eigenvalues of the adjacency matrix. The cones that correspond to the clique numbers for quantum graphs can be characterized. An investigation is needed to determine which Motzkin-Straus clique numbers the Mycielski transformation preserves.
@@ 5.16. Others
Given $k$ lists $L_{1},\cdots,L_{k}$, each consisting of uniformly random $n$-bit strings of size $2^{\frac{n}{k}}$, the $k\text{-}xor$ problem with lists aims to find a $k$-tuple $x_{1},\cdots,x_{k}\in L_{1}\crossproduct\cdots\crossproduct L_{k}$ such that $x_{1}\oplus\cdots\oplus x_{k}=0$, if such a tuple exists.
In the Quantum-Accessible Quantum Memory (QAQM) model of quantum computation, the quantum computation can use as many qubits as needed. The question is whether there are better algorithms to solve the $k\text{-}xor$ problem without QAQM or even in the low-qubits model.
@@
The problem of maximum cardinality matching can be optimized using 2000Q, a QA hardware from D-Wave. A better version of the QUBO formulation was used for this purpose. As part of their energy gap scale investigations, graphs with up to $22$ edges were fabricated. It is worth exploring whether this analysis can be extended to a $G_{2}$ graph. Doing so would provide direct insights into the original Graph family and enable verification of its scaling behavior with the maximum match problem for the $G_{n}$ graph family.
@@
The Measurement-Based Quantum Computation (MBQC) model is a relevant approach to consider in the field of quantum computation. However, a circuit in the usual circuit model transformed to the MBQC model may result in a considerable increase in the circuit width. This is because while the usual circuit model allows for arbitrary wire gate application, the MBQC model only permits interactions between adjacent wires. Furthermore, the use of long-range gates is costly. In light of this, one may inquire whether a protocol can be suggested for both models. Alternatively, one may focus on developing a protocol specifically tailored to the usual circuit model.
@@
The scholarly work cited in (anshu2023nlts, ) posits that quantum local testability is not a necessary condition for the construction of the work. The property, as established in the aforementioned work (Property 1), suffices for clustering the classical distributions of low-energy states. However, it is important to note that the property in question is weaker than local testability. Extant literature establishes that local testability implies NLTS. It is worthwhile to investigate the implications codes with Property 1 have for the quantum PCP conjecture (aharonov2013guest, ).
@@
The study discussed in (anshu2023nlts, ) has successfully established a property referred to as Property 1. Upon analyzing from the perspective of chain complexes, Property 1 seems to exhibit a relationship with the small-set boundary and co-boundary expansion as specified under Definition 1.2 of (hopkins2022explicit, ). This relationship may introduce a classical analog of the NLTS property. Furthermore, (hopkins2022explicit, ) presents a construction technique for classical Hamiltonians that are hard to approximate. The construction technique leverages the sum-of-squares hierarchy.
@@
An extant approach has been developed to tackle the maximum independent set (MIS) problem using quantum wires. Specifically, this method offers a solution for non-planar graphs and high-dimensional ones. Quantum wires, a set of auxiliary wire atoms, have been proposed to facilitate strong interactions between distant atoms (kim2022rydberg, ), thus allowing the synthesis of an entire target graph, $G_{T}$, from a simple first graph, $G_{0}$. However, the applicability of quantum wires in other optimization problems beyond MIS remains uncertain.
@@
The minimum and maximum weight matching problems can be solved in $\mathcal{O}(n(m+n\log n))$ time. There exists an algorithm that finds a maximum weight augmenting path in graph $G$ in $\mathcal{O}(m+n\log n)$ time. The maximum weight matching can be found by enlarging the matching with the path in $G$ in $n$ steps. It is unclear if a quantum algorithm exists for the minimum and maximum weight matching, but if there is, it could improve the running time of the classical algorithm (dorn2009quantum, ).
@@
Is it possible to add iterative parameter tuning and new problem classes, specifically the challenging and low-precision Maximum Independent Set (MIS) problem, to the existing method of generating arbitrary structured problems for D-Wave quantum processing units (QPUs) (8477865, )? This can help compare QPU performance with classical algorithms and identify potential use cases for quantum annealing in real-world applications.